A faulty car causing obstruction at the center of a road. And when a force of 1,500N was applied to push it to a near by mechanic workshop the work done was 10,000J. How far away was the mechanic workshop.
1
Expert's answer
2019-11-19T07:49:30-0500


From the formula for work

"A=F \cdot S \cdot \cos(\alpha)"

Where alpha ("\alpha" )is the angle between the direction of the force and the direction of movement

In our case "\alpha = 0"

then write

"S=\frac{A}{F \cdot \cos(\alpha)}=\frac{10000(J)}{1500(N) \cdot \cos(0)}=6,667(m)"